OpenStudy (chevellemarie01):

Help ASAP!! And explain please Write -1/4x-6 in standard form using integers a)x+4y=-6 b)x-4y=-24 c)x-4y=24 d)x+4y=-24

OpenStudy (chevellemarie01):

@sammixboo can you help me??

OpenStudy (chevellemarie01):

@triciaal

OpenStudy (chevellemarie01):

@sleepyjess

OpenStudy (astrophysics):

Clarify your problem please

OpenStudy (astrophysics):

\[-\frac{ 1 }{ 4 }x-6\]

OpenStudy (chevellemarie01):

thats correct yes

OpenStudy (astrophysics):

Standard form is represented as such \[ax+by=c\]

OpenStudy (chevellemarie01):

okay

OpenStudy (astrophysics):

Is your problem suppose to be \[-\frac{ 1 }{ 4 }x-6 = y?\]

OpenStudy (chevellemarie01):

im not sure i posted the question exactly how my assignment says

OpenStudy (astrophysics):

What you gave wasn't an equation, check your problem again

OpenStudy (chevellemarie01):

its supposed to be y=-1/4-6

OpenStudy (chevellemarie01):

sorry you were right, i typed it wrong the first time

OpenStudy (astrophysics):

So to put it in standard form add 1/4x on both sides

OpenStudy (astrophysics):

what happens

OpenStudy (chevellemarie01):

how do i put it in standard form

OpenStudy (astrophysics):

Standard form is \[ax+by=c\] where a,b, and c are constants, so all you have to do is add 1/4x on both sides and it will cancel it out on the right and then you'll have it in standard form.

OpenStudy (astrophysics):

Constants meaning any number

OpenStudy (chevellemarie01):

Im so lost

OpenStudy (astrophysics):

\[y=-\frac{ 1 }{ 4 }x-6\] how do we move -1/4x on the left side?

OpenStudy (chevellemarie01):

Add it on both sides?

OpenStudy (astrophysics):

Exactly, so what do you get once you do that?

OpenStudy (chevellemarie01):

x+4y=-6, right?

OpenStudy (astrophysics):

No,\[\color{red}{+\frac{ 1 }{ 4 }x}+ y=-\frac{ 1 }{ 4 }x-6\color{red}{+\frac{ 1 }{ 4 }x}\]

OpenStudy (astrophysics):

So it gets cancelled on the right side and we have \[\frac{ 1 }{ 4 }x+y=-6\] now we have to put it with integers, meaning we can't have fractions

OpenStudy (astrophysics):

So you can multiply through by 4

OpenStudy (chevellemarie01):

so we would have x+4y=-24

OpenStudy (astrophysics):

Yes

OpenStudy (chevellemarie01):

thank you so much

OpenStudy (chevellemarie01):

do you think you could help me with another problem?